OpenStudy (anonymous):

An angle bisector of a triangle divides the opposite side of the triangle into segments 6 cm and 5 cm long. A second side of the triangle is 6.9 cm long. Find the longest and shortest possible lengths of the third side of the triangle.

OpenStudy (anonymous):

Do you have its illustrated picture, Ducky?

OpenStudy (anonymous):

I'm afraid not; the text is all we have to go on.

OpenStudy (anonymous):

It's not right triangle?

OpenStudy (anonymous):

No, the triangle type isn't specified.

OpenStudy (anonymous):

Many thanks!

OpenStudy (anonymous):

Because the bisector line in this case ISN'T perpendicular with base, apply golden rule: a/5 = ( a + 5)/ a ->a² = 5 (a +5) ->a² - 5a -25 = 0 => a = (+5 ± √125 )/ 2 = 8.1